OpenStudy (anonymous):

If 'm' is the greatest prime number of 38 and 'n' is the greatest prime number of 100, what is the value of 'm' and 'n.'

OpenStudy (anonymous):

do u mean prime factor?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

ok as 100 = 2*2*5*5 and 38 = 2*19 that makes m = 19 and n = 5